Suppose perfect competitive firm short run cost function



total cost=1/3q3+3q2+10Q+40



. if the market price of the commodity is birr 26 per unit



A, determine the profit maximizing level of out put



B find average fixed cost ,average cost ,average variable cost and marginal cost of firm at optimum level of out put



C find maximum profit of the firm

1
Expert's answer
2021-12-05T18:56:38-0500

"Given\\

Total\space Cost = \frac{1}{3}q^3+3q^2+10Q+40\\

Market \space Price = 26"


a.

Profit Maximizing output is archived where MR = MC

Now Calculate MC

"MC = \frac{∂TC}{∂q}\\MC = q^2 + 6q + 10"

Now put the value in the condition where MR = MC

"q^2 + 6q + 10 = 26\\q^2 + 6q − 16 = 0\\(q−2)(q+8) = 0\\q = 2 \space and\space −8"

quantity should be always positive therefore q = 2

hence when q = 2, profit should be maximized.

b.

"Total\space Cost = \frac{1}{3}q^3+3q^2+10Q+40\\"

Calculation of Average Fixed Cost

"AFC = \frac{TFC}{q}\\AFC = \frac{40}{q}\\where\space q = 2\\AFC =\frac{40}{2} = 20"

Calculation of Average Cost

"AC = \frac{TC}{q}\\AC = \frac{\frac{1}{3}q^3+3q^2+10q+40}{q}\\AC = \frac{2.67 + 12 +20+40}{2}\\AC = 37.335"

Calculation of Average Cost

"AVC = \frac{TVC}{q}\\AVC = \frac{\frac{1}{3}q^3+3q^2+10q}{q}\\AVC =\frac{ 2.67 + 12 +20}{2}\\AVC = 17.335"

Calculation of MC

"MC = q^2 + 6q + 10\\MC = 22 + 6\times 2 + 10\\MC = 4 + 12 + 10\\MC = 26"

c.

"Profit = TR - TC\\

Profit = 26\times 2 - \frac{1}{3}\times 2^3+3\times 2^2+10\times2+40\\

Profit = 112 - 74.67\\

Profit = 37.33"